OpenStudy (anonymous):

The population in the South in 2000 was 100,236,820. In 2010, the population in the South had increased by 14.3% from 2000. About many people lived in the South in 2010?

OpenStudy (anonymous):

C = 100*(y-x)/x. Here, we are given x (the population in 2000), and we need to solve for y (the population in 2010). Solving for y gives y = Cx/100 + x. Substitute the given numbers (C = 14.3, x = 100,236,820): y = 14.3(100,236,820)/100 + 100,236,820 = 114,570,685 (to the nearest whole person).

OpenStudy (anonymous):

do you understand @lozanod684

OpenStudy (anonymous):

Yes Thank you

OpenStudy (anonymous):

best response?

OpenStudy (anonymous):

yes